OpenStudy (anonymous):

Please select the correct slope of the line through the points (-2, 3) and (1, 4)

OpenStudy (luigi0210):

Do you know the slope formula?

OpenStudy (anonymous):

Where are the options?

OpenStudy (texaschic101):

slope formula is : slope(m) = (y2 - y1) / (x2 - x1) (-2,3) x1 = -2 and y1 = 3 (1,4) x2 = 1 and y2 = 4 sub your info into the formula m = (4 - 3) / (1 - (-2) can you solve it from here ?